# MuseGraph: Graph-oriented Instruction Tuning of Large Language Models for Generic Graph Mining

### 02 Mar 2024 v1

@@ 4.4. Case Studies (RQ3)
To demonstrate the advantages of MuseGraph over the LLM-based methods on the generation tasks, we provide the abstract of graph-to-text task prompted by LLaMa, GPT-3.5, and our MuseGraph on AGENDA. The provided graph structure and its ground-truth abstract are displayed at the top of Table 6. We standardized the process by employing our compact graph description as the input across all models.
As shown in Table 6, the abstract generated by LLaMA offers a basic summary while overlooking some key details, such as the specific incremental learning aspect of the fuzzy algorithm (shown in the blue text). Compared with LLaMA, Chatgpt-3.5 generates even three times longer content with an elaborate description. However, the model inevitably includes some fabricated sentences, such as the “Through extensive experiment and evaluation, …” in red, which is irrelevant to the provided ground-truth abstract.
Since our method performs graph-aware instruction tuning with compact graph descriptions, we can produce precise abstracts by extracting the key information from the graph. Moreover, by incorporating the distilled reasoning and generative capacity from CoT-based instructions, MuseGraph can not only produce an abstract that aligns with the core details of the ground-truth but also provides fluent and precise CoT inherited the advanced large language models.
@@ A.2. Details of Compared Baselines
  • ChatGPT-3.5 [41] uses Graph Neural Networks (GNNs) and transformer-based models to generate coherent text based on input structured graphs.
@@ A.3. Instruction Templates across Different Tasks and Datasets
Standard Instruction: Given the target PAPER with the compact graph description, which of the following subcategories of computer science does this PAPER belong to {the category list}. Directly give the most likely category of this PAPER.
@@
Input: The compact graph description of this PAPER 1 is listed as follows: Title: {title of the PAPER 1}. Abstract: {abstract of the PAPER 1}. Ego graph nodes: {ego graph node list}. One-hop neighbors: {1-hop neighbor list}. Random walks: {random walk paths}. The compact graph description of this PAPER 2 is listed as follows: Title: {title of the PAPER 2}. Abstract: {abstract of the PAPER 2}. Ego graph nodes: {ego graph node list}. One-hop neighbors: {1-hop neighbor list}. Random walks: {random walk paths}.
Standard Instruction (pair-wise): Given the compact graph description in PAPER 1 and Paper 2. If the connection between the PAPERs represents the relationship between them, are they connected? Give me a direct answer of ”yes” or ”no”.
@@
CoT Instruction: Given the generated abstract of the target PAPER. Please use the provided compact graph description, to examine how these elements influenced the generation of the abstract with a clear Chain of Thought (CoT). Keep the CoT brief and to the point.

### 13 Mar 2024 v2

@@ 4.4. Case Studies (RQ3)
To demonstrate the advantages of MuseGraph over the LLM-based methods on the generation tasks, we provide the abstract of the graph-to-text task prompted by LLaMA, GPT-3.5, and our MuseGraph on AGENDA. The provided graph structure and its ground-truth abstract are displayed at the top of Table 6. We standardize the process by employing our compact graph description as the input across all models.
As shown in Table 6, the abstract generated by LLaMA offers a basic summary while overlooking some key details, such as the specific incremental learning aspect of the fuzzy algorithm (shown in the blue text). Compared with LLaMA, GPT-3.5 generates even three times longer content with an elaborate description. However, the model inevitably includes some fabricated sentences, such as the “Through extensive experiment and evaluation, …” in red, which is irrelevant to the provided ground-truth abstract.
Since our method performs graph-aware instruction tuning with compact graph descriptions, we can produce precise abstracts by extracting the key information from the graph. Moreover, by incorporating the distilled reasoning and generative capacity from CoT-based instructions, MuseGraph can not only produce an abstract that aligns with the core details of the ground truth but also provides fluent and precise CoT inherited from the advanced large language models.
@@ A.2. Details of Compared Baselines
  • GPT-3.5 [41] uses both Graph Neural Networks (GNNs) and transformer-based models to generate coherent text based on input structured graphs.
@@ A.3. Instruction Templates across Different Tasks and Datasets
Standard Instruction: Given the target PAPER with the compact graph description in the Arxiv dataset, which of the following subcategories of computer science does this PAPER belong to {the category list}. Directly give the most likely category of this PAPER.
@@
Input: The compact graph description of this PAPER 1 is listed as follows: Title: {title of the PAPER 1}. Abstract: {abstract of the PAPER 1}. Ego graph nodes: {ego graph node list of the PAPER 1}. One-hop neighbors: {1-hop neighbor list of the PAPER 1}. Random walks: {random walk paths of the PAPER 1}. The compact graph description of this PAPER 2 is listed as follows: Title: {title of the PAPER 2}. Abstract: {abstract of the PAPER 2}. Ego graph nodes: {ego graph node list of the PAPER 2}. One-hop neighbors: {1-hop neighbor list of the PAPER 2}. Random walks: {random walk paths of the PAPER 2}.
Standard Instruction (pair-wise): Given the compact graph descriptions of PAPER 1 and PAPER 2 in the Arxiv dataset. If the connection between the PAPERs represents the relationship between them, are they connected? Give me a direct answer of “yes” or “no”.
@@
CoT Instruction: Given the generated abstract of the target PAPER in the AGENDA dataset. Please use the provided compact graph description, to examine how these elements influenced the generation of the abstract with a clear Chain of Thought (CoT). Keep the CoT brief and to the point.